OpenStudy (anonymous):

find the area of the region enclosed by the graphs of y=4x^2 and y=2x

OpenStudy (anonymous):

You take the integral of the top function y=2x minus that of the bottom function y=4x^2 from 0 to .5

OpenStudy (anonymous):

\[\int\limits_{0}^{0.5}2x-4x ^{2}dx\]

OpenStudy (dumbcow):

Note: \[\int\limits_{}^{}x^{n} dx = \frac{x^{n+1}}{n+1}\]